2.-(50) End Is, ICVCE and the region of operation, 20012 IM AW 3v 62 200
Answer #1

Solution Given Circuit IE (M. 2001 + B=200 3v YIC KUL equation IB 3- 200 IE 106 113 -0.7=0 Assuming it is Active region -) 2-Voltage at nodes, E, B and c Vc = 3 - Pe x200 = 3 - 200x 444-21 x10-6 Vc = 2-911 volt VE こOV VB = Hence VBE =Ve- Imax Ig = 2-